OpenStudy (anonymous):

The cost of buying 5 slices of pizza and 2 drinks at a pizzeria is $15.50. The cost of buying 6 slices of pizza and 6 drinks is $24. Assume all slices are the same price and all drinks are the same price. What is the price of each? a. Each slice costs $1.50, and each drink costs $2.50. b. Each slice costs $2.00, and each drink costs $2.00. c. Each slice costs $2.50, and each drink costs $1.50. d. Each slice costs $3.00, and each drink costs $0.25.

OpenStudy (anonymous):

@RyleeC

OpenStudy (anonymous):

5p+2d=15.5 6p+6d=24 p=-d+4 5(-d+4)+2d=15.5

OpenStudy (anonymous):

1.5 ; A ?

OpenStudy (anonymous):

Nope, d is for drinks so if d=1.5 its...?

OpenStudy (anonymous):

Oh, Its C

OpenStudy (anonymous):

Yep.